OpenStudy (anonymous):

Please help im so confused Water coming out from a fountain is modeled by the function f(x) = -x^2 + 7x + 5 where f(x) represents the height, in feet, of the water from the fountain at different times x, in seconds. What does the average rate of change of f(x) from x = 4 to x = 6 represent?

OpenStudy (anonymous):

The water falls down with an average speed of 5 feet per second from 4 seconds to 6 seconds. The water falls down with an average speed of 3 feet per second from 4 seconds to 6 seconds. The water travels an average distance of 5 feet from 4 seconds to 6 seconds. The water travels an average distance of 3 feet from 4 seconds to 6 seconds

OpenStudy (imstuck):

The water falls down with an average speed of 3 feet per second from 4 seconds to 6 seconds. That's your answer.

OpenStudy (imstuck):

Wanna know how to find it?

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

and thank you

OpenStudy (imstuck):

k, they give you the function:\[f(x)=-x ^{2}+7x+5\]

OpenStudy (imstuck):

In order to find the y values at these x values, replace the x in the equation with a 4 and then a 6. When you replace the x with 4, you get this:

OpenStudy (imstuck):

\[f(4)=-(4)^{2}=7(4)+5\]

OpenStudy (imstuck):

That gives you a y value of 17. The coordinates of that are (4, 17). That means that when x is 4, y is 17.

OpenStudy (imstuck):

Now let's do the same for f(6).

OpenStudy (imstuck):

\[f(6)=-(6)^{2}+7(6)+5\]When you do that you get a y value of 11.

OpenStudy (imstuck):

so the coordinates of that point are (6, 11), which means that when x is 6, y is 11. Now find the average rate of change between those 2 points. Hint: the average rate of change between two points is also known as the slope. So use the slope formula to find your answer.

OpenStudy (imstuck):

\[\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]and sub in your values:

OpenStudy (anonymous):

so \[\frac{ 11-17 }{ 6-4 }\]

OpenStudy (imstuck):

\[\frac{ 11-17 }{ 6-4 }=\frac{ -6 }{ 2 }=-3\]That means that the rate of change from x = 4 to x = 6 is falling 3 feet per second. Falling because the sign on the 3 is negative.

OpenStudy (anonymous):

oh this helped more than u could know I never understood how to do average rate of change. thank you so much

OpenStudy (imstuck):

You are so very welcome, my dear!